A car accelerates uniformly from rest to 16.0 m/s in 7.24 s along a level stretch of road. Ignoring friction, determine the average power required to accelerate the car if (a) the weight of the car is 6.38 x 10^3 N, and (b) the weight of the car is 1.45 x 10^4 N.

Please do not round answer
Answer #1

Power=work done/time

As we know Work done by force (F=ma) is equal to change in kinetic energy

W=F.d=1/2m(Vf^2 -Vi^2) , Vi=0, Vf=16m/s, g=9.81m/s^2

W1=6.38^10^3 N, W2=14.5*10^3N, t=7.24

a)

P=1/2 *(WV^2/g) /t=(0.5*6.38*10^3 *16*16) /7.24*9.81

= 11.49799*10^3 =11497.99 joule

b)

P=0.5*1.45*10^4 *16*16/7.24*9.81=26131.79 joule